OpenStudy (anonymous):

derivative xe^((-x^2)/2)

OpenStudy (anonymous):

chain rule plus product rule. \[(fg)'=f'g+g'f\] with \[f(x)=x,f'(x)=1,g(x)=e^{-\frac{x^2}{2}},g'(x)=-xe^{-\frac{x^2}{2}}\]

OpenStudy (anonymous):

\[ xe ^{-x ^{2}/2}\] \[(1-x ^{2})\times e ^{-x ^{2}/x}\]

OpenStudy (anonymous):

that "x" in the middle is multiplication